OpenStudy (strawberry17):

(1 - sin theta) ÷ (sin theta) = (cot^2 theta) ÷ (1 + csc theta)

zepdrix (zepdrix):

Hmm that's a strange one :b thinking....

OpenStudy (strawberry17):

wait, the first side, the bottom denominator is just sine theta

zepdrix (zepdrix):

\[\Large \frac{1-\sin \theta}{\sin \theta}\quad=\quad \frac{\cot^2\theta}{1+\csc \theta}\]Oh hah :3 my bad

zepdrix (zepdrix):

Ok ok ok this one isn't too bad. Just a few weird steps to take :o We're going to manipulate the left side, and try to match it to the right. Let's start by multiplying the top and bottom by (1+sin theta)\[\Large \frac{1-\sin \theta}{\sin \theta}\left(\frac{1+\sin \theta}{1+\sin \theta}\right)\quad=\quad \frac{?}{\sin \theta(1+\sin \theta)}\]

zepdrix (zepdrix):

Do you understand what happens in the numerator?

OpenStudy (strawberry17):

Yes! That is the first step that I took also, but could not finish.. :)

zepdrix (zepdrix):

So that puts us here, yes?\[\Large \frac{1-\sin^2\theta}{\sin \theta(1+\sin \theta)}\]Does the top simplify any further?

OpenStudy (strawberry17):

yes

OpenStudy (strawberry17):

to cos^2 theta

zepdrix (zepdrix):

\[\Large \frac{\cos^2\theta}{\sin \theta(1+\sin \theta)}\]Hmm ok good.

zepdrix (zepdrix):

Recall that:\[\Large \cot \theta\quad=\quad \frac{\cos \theta}{\sin \theta}\qquad\implies\qquad \cot^2\theta\quad=\quad \frac{\cos^2\theta}{\sin^2\theta}\]

zepdrix (zepdrix):

Hmm see how we have a cos^2 on top, and a sin on bottom? If we only had one more sine in the bottom, we could write our numerator as cot^2. Hmm we're so close! Where can we get another sine from?

OpenStudy (strawberry17):

distributing the sine on the bottom?

zepdrix (zepdrix):

Hmm we actually want to do the opposite! We want to steal a sine from inside of the brackets. So we'll `factor out` sine from each term in the brackets.

zepdrix (zepdrix):

\[\Large \frac{\cos^2\theta}{\sin \theta(1+\sin \theta)}\quad=\quad \frac{\cos^2\theta}{\sin \theta\cdot \sin \theta\left(\frac{1}{\sin \theta}+1\right)}\]Understand what I did there? :o

OpenStudy (strawberry17):

not really :O

OpenStudy (strawberry17):

where did the 1 over sine come from?

zepdrix (zepdrix):

\[\LARGE\color{royalblue}{1+\sin \theta\qquad=\qquad \sin \theta\left(\frac{1}{\sin \theta}+\frac{\sin \theta}{\sin \theta}\right)}\]I factored a sine out of each term inside the brackets.

zepdrix (zepdrix):

Multiplying the sines together in the denominator gives us\[\Large \frac{\cos^2\theta}{\sin \theta(1+\sin \theta)}\quad=\quad \frac{\cos^2\theta}{\sin ^2\theta\left(\frac{1}{\sin \theta}+1\right)}\]

zepdrix (zepdrix):

Using our identity from earlier:\[\Large\frac{\color{orangered}{\cos^2\theta}}{\color{orangered}{\sin ^2\theta}\left(\frac{1}{\sin \theta}+1\right)}\quad=\quad\frac{\color{red}{\cot^2\theta}}{\left(\frac{1}{\sin \theta}+1\right)}\]Yes?

OpenStudy (strawberry17):

Hey @zepdrix just wanted to let you know I'm still on this question and will be back to finish it, just got really busy after the other day. Sorry i didn't respond :o

OpenStudy (strawberry17):

care to help me on it a bit later? :)

zepdrix (zepdrix):

sure :3

OpenStudy (strawberry17):

thanks!

OpenStudy (strawberry17):

k, I'm back!

OpenStudy (strawberry17):

could you explain though, why when dividing a sine in the parentheses of: sine(1+sine) that sine becomes sine^2 . I just don't understand that part... is it just something I'm just supposed to know?

zepdrix (zepdrix):

From this step?\[\Large \frac{\cos^2\theta}{\sin \theta(1+\sin \theta)}\]Factoring a sine out of each term in the brackets gives us,\[\Large \frac{\cos^2\theta}{\sin \theta\cdot \sin \theta\left(\frac{1}{\sin \theta}+1\right)}\]

zepdrix (zepdrix):

That's where the extra sine is coming from, still confused on that? :o

OpenStudy (strawberry17):

yes from that step :) what does factoring out something mean again??

zepdrix (zepdrix):

Here's a quick recap on factoring. Factoring is 2 steps: ~We divide a value out of each term, ~We multiply by that same value, except we save the multiplication, we don't multiply it out. Example:\[\Large 3x+9\]Step 1 is: ~Divide 3 out of each term,\[\Large \frac{3x}{3}+\frac{9}{3}\]But in order to keep the equation balanced, we have to also multiply by 3.\[\Large 3x+9\quad = \quad 3\left(\frac{3x}{3}+\frac{9}{3}\right)\quad=\quad 3(x+3)\]

OpenStudy (strawberry17):

ookkayy that makes sense! thanks for the example really helped.

zepdrix (zepdrix):

\[\Large 1+\sin \theta\]So in our brackets we're dividing each term by sin theta:\[\Large \frac{1}{\sin \theta}+\frac{\sin \theta}{\sin \theta}\]But we must also multiply by sin theta to keep it balanced, yes?\[\Large \sin \theta\left(\frac{1}{\sin \theta}+\frac{\sin \theta}{\sin \theta}\right)\]

zepdrix (zepdrix):

Ya it's a weird process :) factoring is really important though! make sure you understand it.

OpenStudy (strawberry17):

Yes for sure... I will be sure to remember it now! :)

OpenStudy (strawberry17):

k so i guess that it is it? we will just change the cos^2/sine^2 into cot^2 over (1 + sine) ?

zepdrix (zepdrix):

So you created a cot^2, good. But we also need to find a secant somewhere, don't we?

zepdrix (zepdrix):

I mean csc*

OpenStudy (strawberry17):

yes, nvm that isn't it :0

zepdrix (zepdrix):

\[\Large\frac{\cot^2\theta}{\left(\frac{1}{\sin \theta}+1\right)}\]

zepdrix (zepdrix):

It almost perfectly matches up, if it wasn't for one piece. Hmmm Can you think of any identities that might help? +_+

OpenStudy (strawberry17):

1/sine equals to csc ... : )

OpenStudy (strawberry17):

so we have it! oh that wasn't so bad after all : )

zepdrix (zepdrix):

yayyy \c:/ gj strawberry shortcake!

OpenStudy (strawberry17):

you are too fuuny hah : )

zepdrix (zepdrix):

No I'll admit, that was a pretty nasty problem. Required quite a bit of work to get through :3 Gotta really be comfortable with your identities to get through a problem like that.

OpenStudy (strawberry17):

I am mastering them slowly but hey thanks so much for thee help, greatly appreciated! : )

zepdrix (zepdrix):

np c: